# A Clique Partitioning-Based Algorithm for Graph Compression

### v2

@@ 4 Properties of CPGC
The compressed graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGC preserves the path information of the original graph $G(U,W,E)$.
@@
    Provided in Appendix A.3.
In the following, we determine a bound on the number of edges in the compressed graph obtained by CPGC. This bound tells us how good the compression achieved by CPGC is. First, we state a theorem from [9] that guarantees the existence of a $\delta$-clique in a bipartite graph. This theorem will be used in the proofs of CPGC’s compression properties.
Theorem 4.2
@@
Next, we provide a bound on the minimum number of edges $\hat{m}$ of $G$ required to obtain a compression of $G$ into $G^{*}$ by extracting $\delta$-cliques.
Lemma 4.1
Given a graph $G(U,W,E)$, where $n=|U|=|W|$, $m=|E|$, and a constant $\delta$, $10<\delta<$ 1, if $m<2n^{2-\frac{\delta}{2}}$ then extracting $\delta$-cliques and replacing them with tripartite graphs as done in CPGC does not lead to a compression of $G$.
@@
    Provided in Appendix A.3.
Theorem 4.3
Let $G(U,W,E)$ be any bipartite graph with $|U|=|W|=n$ and $|E|=m>2n^{2-\frac{\delta}{2}}$, where $\delta$ is a constant such that $10<\delta<$ 1. Then, the number of edges in the compressed graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGC is $|E^{*}|=O\left(\frac{m}{k(n,m,\delta)}\right)$.
@@
    We follow the basic idea of the proof from Theorem 2.4 in [9] and extend it to apply to the CPGC algorithm. We assume that initially the compressed graph $G^{*}$ has 0 edges, i.e., $|E^{*}|=0$, and that edges are added to $E^{*}$ as the algorithm progress. To estimate the number of edges in $E^{*}$, we divide the iterations of CPGC into stages, where each stage $i$ consists of extracting one or more $\delta$-cliques with a fixed $\hat{k}_{i}$. Therefore, the $i^{th}$ stage includes all $\delta$-cliques extracted after the number of edges in $G$ becomes less than $m/2^{i-1}$ for the first time, and before the number of edges in $G$ becomes less than $m/2^{i}$ for the first time. For stage $i$, $\hat{k}$ is always going to be at least $k_{i}=\frac{\delta\log n}{2\log(2^{i}\cdot 2n^{2}/m)}$.
@@
    CPGC terminates extracting cliques when the number of remaining edges in $G$ is less than $2n^{2-\frac{\delta}{2}}$. This is to eliminate the extraction of trivial cliques, as shown in Lemma 4.1. Next, we determine an upper bound on the total number of edges $m^{*}$ added by CPGC to $G^{*}$ before it terminates extracting $\delta$-cliques (i.e., when $m<2n^{2-\frac{\delta}{2}}$). Therefore,
    (4.1) $\begin{split}m^{*}\leq\sum_{i=1}^{\big{\lceil}\log\left(\frac{m}{2n^{2-\delta/2}}\right)\big{\rceil}}\bigg{(}\frac{1}{k_{i}}+\frac{1}{n^{1-\delta}}\bigg{)}\frac{2m}{2^{i}}\\ \leq\sum_{i=1}^{\infty}\bigg{(}\frac{1}{k_{i}}+\frac{1}{n^{1-\delta}}\bigg{)}\frac{2m}{2^{i}}.\end{split}$
@@
(4.2) $\begin{split}m^{*}\leq\frac{4m}{k}\sum_{i=1}^{\infty}\frac{1}{2^{i}}+\frac{4m}{\delta\log n}\sum_{i=1}^{\infty}\frac{i}{2^{i}}+\frac{2m}{n^{1-\delta}}\sum_{i=1}^{\infty}\frac{1}{2^{i}}\\ \leq 2m\left(\frac{2}{k}+\frac{4}{\delta\log n}+\frac{1}{n^{1-\delta}}\right).\end{split}$
Thus, $m^{*}=O\left(\frac{m}{k}\right)$. After CPGC finishes extracting cliques, there are $2n^{2-\delta/2}$ edges remaining in $G$. Those remaining edges are trivial cliques (i.e., single edges) that are added to $G^{*}$. The number of remaining edges $2n^{2-\delta/2}$ is in $O\left(\frac{m}{k}\right)$. Therefore, the number of edges in the compressed graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGC is $|E^{*}|=O\left(\frac{m}{k}\right)$.
Extension to non-bipartite graphs. In Appendix A.4 we describe how CPGC can be extended to compress non-bipartite graphs.
@@ 5 Experimental Results
We investigated the performance of CPGC in terms of both running time and compression ratio, and compared it with the FM algorithm proposed by Feder and Motwani [9]. We could not locate any existing implementation of the FM algorithm. Thus, we independently implemented it in C to the best of our ability. However, we encountered a limitation with the FM algorithm regarding large graphs, where the calculations for selecting vertices may exceed the machine representation limits. Consequently, we could only obtain results for bipartite graphs with up to $n=128$ vertices in each bipartition and 16 thousand edges (i.e., small graphs). To ensure a fair comparison between algorithms, we excluded the time spent on extracting trivial cliques ($\hat{k}=1$) when measuring the runtime of the FM algorithm. CPGC addresses this by excluding trivial cliques considering $\hat{k}>1$, as indicated in Line 6 of Algorithm 1. Furthermore, we investigated the performance of our algorithm, CPGC, on large bipartite graphs with up to 32 thousand vertices in each bipartition and approximately 1.05 billion edges. Both CPGC and FM algorithms were implemented in C and experiments were conducted on a Linux system with an AMD EPYC-74F3 processor, 1 CPU core, 3.2GHz, and 128 GB of memory. We used the GCC compiler (version 8.5.0) for compiling and executing the C code.
Figure 3: CPGC: Average compression ratio (top row: (a), (b), and (c)) and running time (bottom row: (d), (e), and (f)) for graphs with 3.36 million to 1.05 billion edges.
For our experiments, we generated bipartite graphs with various densities using the $G(n,p)$ random graph model adapted to bipartite graphs. In this model, an edge is included in the bipartite graph with probability $p$, independently from every other edge. We implemented a Python program to generate instances of bipartite graphs with the number of vertices $|U|=|W|=n$, where $n$ ranges from 32 to 32,768, and densities of 0.8, 0.85, 0.90, 0.95, and 0.98. This generated graphs with up to 1.05 billion edges, considering $p$ as a measure of the density of the generated bipartite graphs. For each bipartite graph with a given density, we created 10 different instances. We then ran each instance for six different values of delta ($\delta$: 0.5, 0.6, 0.7, 0.8, 0.9, 1), presenting the average values and standard deviations of the running time and compression ratio for these 10 instances.
5.1 Results for Large Graphs.
We present the results obtained from testing CPGC on large bipartite graphs, where each bipartition consists of $n=2^{i}$ vertices, with $i$ ranging from 11 to 15, and a number of edges $m$ ranging from approximately 3.36 million to 1.05 billion. Figures 3a-3c show the average compression ratio, while Figures 3d-3f show the average running time for bipartite graphs.
Average compression ratio. The compression ratio for CPGC is calculated as $\frac{m}{m^{*}}$, where $m$ and $m^{*}$ represent the number of edges in the given and compressed graphs, respectively. We observe that, for the same $n$ and $\delta$, the compression ratio increases with density. For instance, with $n=2^{15}$ and $\delta=0.5$, at a density of $0.80$ with $m=214.75$ million, the compression ratio is $1.95$. In contrast, at a density of $0.98$, with $m=1.05$ billion, the compression ratio increases to $2.44$. This is because the probability of finding common neighbours is correlated with the density, that is, increasing the density increases the probability of finding a large set of common neighbours of the right partition of the $\delta$-clique. Finding more common neighbours results in CPGC extracting more edges, and thus increasing the compression ratio. However, when the density is kept constant, the compression ratio is not monotonic and depends on $\delta$ and $n$. For instance in Figure 3a, when $n=2^{13}$ and $m=53.69$ million for density $0.80$, the compression ratio initially increases from 1.85, for $\delta=0.5$, to 2, for $\delta=0.6$, achieving the maximum compression. Yet, further increasing $\delta$ to 1 leads to a reduction in the compression ratio to 1.36. This is because increasing $\delta$ increases the number of vertices in the right partition $K_{q}$ of a $\delta$-clique, as $|K_{q}|=\hat{k}$ and $\hat{k}\propto\delta$. Consequently, this decreases the probability of finding the common neighbors, and therefore for lower density, increasing $\delta$ results in a decrease in the compression ratio. Conversely, as the density increases, the probability of discovering common neighbors also increases. Hence, at a density of 0.98, an increase in $\delta$ correlates with an increase in the compression ratio for all $n$ and $m$ values (Figure 3c). It is important to note that the compression ratio depends on $n$, $m$, and $\delta$. Certain combinations of $n$, $m$, and $\delta$ lead to the highest compression ratio. We leave the investigation of finding the optimal combination of these parameters for future work.
@@
Figure 4: CPGC vs. FM: Compression ratio relative to FM (top row: (a), (b) and (c)) and speedup relative to FM (bottom row: (d), (e) and (f)) for graphs with 819 to 16 thousand edges.
Average running time. We observe that for the same $\delta$ and density, the running time increases as the number of vertices increases which is expected (CPGC running time is $O(mn^{\delta})$). We also observe that for the same $n$ and density, the running time decreases as $\delta$ increases. For instance, with $m=214.75$ million for density = 0.80 and $\delta=0.5$, the running time is 176.27 seconds, whereas for $\delta=1$ it reduces to 117.8 seconds. This is because of the relation between $\delta$ and $\hat{k}$ described earlier, CPGC cannot find a large set of common neighbours which results in a lower compression ratio for lower density. Also, when CPGC does not find an edge with $u_{i}\in U$ for any vertex $w_{j}\in K_{q}$, then it skips to the next vertex in $U$ to check for an edge with $w_{j}\in K_{q}$ thus, decreasing the running time. Therefore, we do not observe an increase in the running time as $\delta$ increases. However, for density = 0.98, the running time decreases as $\delta$ increases, even when the compression ratio increases. For example, with $m=1.05$ billion for density = 0.80 and $\delta=0.5$, the running time is 164.83 seconds, whereas for $\delta=1$ the running time is 150.97 seconds. This is because for higher density and large $\delta$, large $\delta$-cliques are extracted which decreases the total number of iterations performed by CPGC and subsequently reduces the running time.
Figure 5: Dinitz($G$) vs. Dinitz($G^{*}$): Average running time of Dinitz’s algorithm on the original bipartite graph $G$ (labeled Dinitz($G$)) and on the compressed graph $G^{*}$ (labeled Dinitz($G^{*}$)), for a large graph with approximately 32,000 vertices in each bipartition and 214.75 million to 1.05 billion edges corresponding to densities = $0.80,0.90,$ and $0.98$ and different $\delta$.
5.2 Results for Small Graphs.
@@
In this section, we compare the performance of CPGC against FM on small graphs, particularly, on graphs with $n=2^{i}$ vertices in each bipartition, where $i$ ranges from 5 to 7 and the number of edges ranging from 819 to 16,056. However, due to the limitations of FM as previously mentioned, our comparison is constrained to $n=2^{7}$ and $m=16,056$. We compare the performance of the two algorithms using two metrics, relative compression ratio, and relative speedup.
Compression ratio relative to FM. We define the compression ratio relative to FM, as the ratio of compression ratio achieved by CPGC and FM. Figures 4a-4c show the relative compression ratio relative to FM achieved by CPGC in the case of small graphs, for $\delta=0.5,0.6,0.7,0.8,0.9,1$, and densities = 0.80, 0.90, 0.98. In all cases, the relative compression ratio ranges from $1$ to $1.27$, highlighting that the compression ratio achieved by CPGC is at least as that of FM. CPGC achieves a greater compression ratio than FM because it extracts multiple $\delta$-cliques before $\hat{k}$ decreases as shown in Figure 2(b), resulting in removing more edges at higher values of $\hat{k}$ (Section 3.2). We observe that CPGC achieves a compression ratio equal to that of FM for a small graph with $n=32$, $m=819$, density = 0.80, and $\delta=0.5$. This is because a small graph does not require many iterations for extracting $\delta$-cliques, thereby not giving CPGC the opportunity to extract more cliques over the FM algorithm. However, for the same $n$ and $\delta$ when the graph density increases the compression ratio achieved by CPGC also increases. For example, with $n=32$, $m=921$, and $1,003$, the relative compression ratio is 1.03 and 1.25, respectively. As explained before, certain combinations of $n$, $m$, and $\delta$ lead to the highest compression ratio. For instance, when $n=128$, $m=16,056$ for a density of 0.98, we observe that the relative compression ratio initially increases from 1.1 to 1.14 with increasing $\delta$ from 0.5 to 0.7. However, further increasing $\delta$ to 1 results in a decrease in the relative compression ratio to 1.1.

### v3

@@ 4 Properties of CPGR
The restructured graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGR preserves the path information (i.e., vertex-to-vertex reachability) of the original graph $G(U,W,E)$.
@@
Let’s assume that CPGR extracts only one $\delta$-clique $C_{q}$ from the given graph $G$. In the $\delta$-clique $C_{q}$, the right partition $K_{q}\subseteq W$ is formed by selecting $\hat{k}$ vertices from $W$ and the left partition $U_{q}\subseteq U$ is formed with the common neighbors of $K_{q}$, forming a complete bipartite graph with edge set $E_{q}$. The restructured graph $G^{*}$ is formed with the same left and right partitions $U$ and $W$ of the graph $G$, and a third partition $Z$ which is a set of additional vertices associated with each of the $\delta$-cliques that CPGR extracts. Our main concern is with the edges $(u_{i},w_{j})\in E^{*}$ in the restructured graph $G^{*}$. $E^{*}$ contains the edges that replace the $\delta$-clique $C_{q}$ by adding the additional vertex $z_{q}\in Z$, and the set of edges $\hat{E}=E\setminus E_{q}$ which were not part of the $\delta$-clique $C_{q}$. It can be easily seen that $\hat{E}\subset E$, that is, the edges in $\hat{E}$ are edges in $G$, and the remaining edges $E^{*}\setminus\hat{E}$ in $E^{*}$ are the edges that replace $C_{q}$ in $G^{*}$. Each edge $(u_{i},w_{j})\in E_{q}$, is replaced by two edges, $(u_{i},z_{q})$ and $(z_{q},w_{j})$, where $u_{i}\in U$, $w_{j}\in W$, and $z_{q}\in Z$. Therefore, for each edge $(u_{i},w_{j})\in E_{q}$, there exists a path from $u_{i}$ to $w_{j}$ composed of two edges, $(u_{i},z_{q})$ and $(z_{q},w_{j})$, that passes through the additional vertex $z_{q}\in Z$, thus preserving the path information. This holds true for all the $\delta$-cliques extracted by CPGR. ∎
In the following, we determine an upper bound on the number of edges in the restructured graph obtained by CPGR. First, we state a theorem from [14] that guarantees the existence of a $\delta$-clique in a bipartite graph. This theorem is necessary in the proof of the bound.
Theorem 2.
@@
Lemma 1.
Given a graph $G(U,W,E)$, where $n=|U|=|W|$, $m=|E|$, and a constant $\delta$, $0<\delta<1$, if $m<2n^{2-\frac{\delta}{2}}$ then extracting $\delta$-cliques and replacing them with tripartite graphs as done in CPGR does not lead to a reduction in the number of edges of $G$.
@@
In the $\delta$-clique based graph restructuring performed by CPGR, the restructured graph $G^{*}(U,W,Z,E^{*})$ is obtained from $G$ by replacing $|E_{q}|=|U_{q}\times K_{q}|$ edges in a $\delta$-clique $C_{q}$ with $|E^{*}|=|U_{q}|+|K_{q}|$ edges in $G^{*}$ and adding an additional vertex $z_{q}\in Z$ for each extracted $\delta$-clique $C_{q}$. The size of the right partition of $C_{q}$ is $|K_{q}|={k}=k(n,m,\delta)=\Big\lfloor{\frac{\delta\log n}{\log(2n^{2}/m)}}\Big\rfloor$. If $m<2n^{2-\frac{\delta}{2}}$ then the size of the right partition $|K_{q}|$ is $k=1$. Therefore, the number of edges in the $\delta$-clique is $|E_{q}|=|U_{q}|\cdot 1=|U_{q}|$. Those edges are replaced by $|E^{*}|=|U_{q}|+\hat{k}=|U_{q}|+1$ edges in $G^{*}$. Thus, replacing the $\delta$-clique in $G$ with $m<2n^{2-\frac{\delta}{2}}$, the number of edges in $G^{*}$ would actually increase by 1, which does not lead to a restructuring of $G$. ∎
Theorem 3.
Let $G(U,W,E)$ be any bipartite graph with $|U|=|W|=n$ and $|E|=m>2n^{2-\frac{\delta}{2}}$, where $\delta$ is a constant such that $0<\delta<1$. Then, the number of edges in the restructured graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGR is $|E^{*}|=O\left(\frac{m}{k(n,m,\delta)}\right)$.
@@
We follow the basic idea of the proof from Theorem 2.4 in [14] and extend it to apply to the CPGR algorithm. We assume that initially the restructured graph $G^{*}$ has 0 edges, i.e., $|E^{*}|=0$, and that edges are added to $E^{*}$ as the algorithm progress. To estimate the number of edges in $E^{*}$, we divide the iterations of CPGR into stages, where each stage $i$ consists of extracting one or more $\delta$-cliques with a fixed $\hat{k}_{i}$. Therefore, the $i^{th}$ stage includes all $\delta$-cliques extracted after the number of edges in $G$ becomes less than $m/2^{i-1}$ for the first time, and before the number of edges in $G$ becomes less than $m/2^{i}$ for the first time. For stage $i$, $\hat{k}$ is always going to be at least $k_{i}=\frac{\delta\log n}{2\log(2^{i}\cdot 2n^{2}/m)}$.
@@
CPGR terminates extracting cliques when the number of remaining edges in $G$ is less than $2n^{2-\frac{\delta}{2}}$. This is to eliminate the extraction of trivial cliques, as shown in Lemma 1. Next, we determine an upper bound on the total number of edges $m^{*}$ added by CPGR to $G^{*}$ before it terminates extracting $\delta$-cliques (i.e., when $m<2n^{2-\frac{\delta}{2}}$). Therefore,
$\begin{split}m^{*}\leq\sum_{i=1}^{\big\lceil\log\left(\frac{m}{2n^{2-\delta/2}}\right)\big\rceil}\bigg(\frac{1}{k_{i}}+\frac{1}{n^{1-\delta}}\bigg)\frac{2m}{2^{i}}\leq\sum_{i=1}^{\infty}\bigg(\frac{1}{k_{i}}+\frac{1}{n^{1-\delta}}\bigg)\frac{2m}{2^{i}}.\end{split}$ (1)
@@
$\begin{split}m^{*}\leq\frac{4m}{k}\sum_{i=1}^{\infty}\frac{1}{2^{i}}+\frac{4m}{\delta\log n}\sum_{i=1}^{\infty}\frac{i}{2^{i}}+\frac{2m}{n^{1-\delta}}\sum_{i=1}^{\infty}\frac{1}{2^{i}}\leq 2m\left(\frac{2}{k}+\frac{4}{\delta\log n}+\frac{1}{n^{1-\delta}}\right).\end{split}$ (2)
Thus, $m^{*}=O\left(\frac{m}{k}\right)$. After CPGR finishes extracting cliques, there are $2n^{2-\delta/2}$ edges remaining in $G$. Those remaining edges are trivial cliques (i.e., single edges) that are added to $G^{*}$. The number of remaining edges $2n^{2-\delta/2}$ is in $O\left(\frac{m}{k}\right)$. Therefore, the number of edges in the restructured graph $G^{*}(U,W,Z,E^{*})$ obtained by CPGR is $|E^{*}|=O\left(\frac{m}{k}\right)$. ∎
Extension to non-bipartite graphs. CPGR can be extended to restructure non-bipartite graphs using a similar technique to that described in [14]. Consider a directed graph $H(V,E)$ with $n$ vertices and $m$ edges. Initially, we transform $H$ into a bipartite graph $G\left(L,R,E^{\prime}\right)$, where each vertex $v\in V$ is duplicated into $v_{L}\in L$ and $v_{R}\in R$. For each directed edge $(u,v)$ in $E$, we create a directed edge $\left(u_{L},v_{R}\right)$ in $E^{\prime}$, with the direction from $L$ to $R$. Furthermore, for each vertex $v\in V$ we add a directed edge from $v_{R}$ to $v_{L}$. The proposed graph restructuring algorithm is then applied to the bipartite graph $G$, which includes only the edges from $L$ to $R$. Once the restructured graph is computed we add to it the $n$ directed edges from $u_{R}$ to $u_{L}$, corresponding to each vertex $u\in V$. Since the path information of the original graph is preserved by this transformation, it allows different graph algorithms to work on the restructured graph as well. Undirected general graphs can be transformed first into directed graphs by simply replacing each undirected edge $(u,v)$ in the original graph by two directed edges $(u,v)$ and $(v,u)$, and then applying the technique presented above to transform the directed graph into a bipartite graph. The time complexity and the compression ratio of the proposed algorithms are the same as in the case of the bipartite graphs. The time complexity and the reduced number of edges of CPGR are the same as in the case of the bipartite graphs. A similar approach for converting a general graph to a bipartite graph and then restructuring the obtained bipartite graph was employed in [3, 10].
@@ 5 Experimental Results
We evaluated the running time and the reduction in the number of edges, $R$ of CPGR by comparing it with the FM algorithm [14]. Let $m=|E|$ be the number of edges in the given graph $G$ and $m^{*}=|E^{*}|$ be the number of edges in the restructured graph $G^{*}$ and $m^{*}\leq m$, then we define the reduction in the number of edge as $R=m-m^{*}$. Due to the lack of an existing implementation, we implemented FM in C. However, its vertex selection calculations encountered machine representation limits on large graphs, restricting our comparative analysis to small bipartite graphs (up to $n=128$ vertices per partition, $\sim$16k edges). To ensure a fair runtime comparison, we excluded trivial clique extraction ($\hat{k}=1$) for both algorithms (CPGR inherently does this as per Algorithm 1, Line 10). We further assessed CPGR on large bipartite graphs (up to 1.05 billion edges) and real-world datasets from the SuiteSparse Matrix Collection [9]. Additionally, we tested the utility of our restructured graph by applying the Dinitz algorithm for maximum bipartite matching to large bipartite graphs and the APSP algorithm on real-world datasets. All algorithms, were implemented in C and tested on a Linux system (AMD EPYC-74F3, 3.2GHz, 1 core, 128 GB RAM) using GCC 8.5.0. The implementation is available at GitHub [17].
Our experiments involved synthetic bipartite graphs generated using adapted $G(n,p)$ random graph model with $|U|=|W|=n$, $n\in[32,32768]$, densities $p\in[0.8,0.98]$, up to 1.05 billion edges and general graphs with standard $G(n,p)$ with $n=32000$, densities $p\in[0.4,0.7]$, $\sim$409 to 716 million edges, with 10 instances per configuration. We also used real-world sparse graphs from SuiteSparse [9]. For each instance, we ran CPGR with six $\delta$ values [0.5 to 1] and show average results and standard deviations for runtime and reduction in number of edges $R$, as well as the performance of the application algorithms on the restructured graphs.
5.1 Results for Small Bipartite Graphs
@@
Figure 3: CPGR vs. FM: Reduction in number of edges relative to FM (top row: (a), (b) and (c)) and speedup relative to FM (bottom row: (d), (e) and (f)) for graphs with 819 to 16 thousand edges.
@@
In this section, we compare the performance of CPGR against FM on small bipartite graphs, particularly, with $n=2^{i}$ vertices in each bipartition, where $i$ ranges from 5 to 7 and the number of edges ranging from 819 to 16,056. However, due to the limitations of FM as previously mentioned, our comparison is constrained to $n=2^{7}$ and $m=16,056$. We compare the performance of the two algorithms using two metrics, speedup, and $R$ relative to FM.
Reduction in number of edges $R$ relative to FM. We define the edge reduction ratio relative to FM as the ratio of the $R$ achieved by CPGR to that of FM. Figures 3a-3c illustrate the relative $R$ of CPGR compared to FM for small bipartite graphs across different $\delta$ values (0.5 to 1) and densities (0.80, 0.90, 0.98). In all these cases, the relative $R$ ranges from 1 to 1.27, indicating that CPGR achieves at least the same, and often a greater reduction than FM. This improved performance of CPGR stems from its ability to extract multiple $\delta$-cliques before $\hat{k}$ decreases (as shown in Figure 2b), allowing for the removal of more edges at higher $\hat{k}$ values (Section 3.2). Notably, for a graph with $n=32$, $m=819$, a density of 0.80, and $\delta$=0.5, CPGR achieves an $R$ equivalent to FM. This is because the small size of the graph limits the number of iterations required for $\delta$-clique extraction, thus not providing CPGR with a significant advantage over FM. However, for the same $n=32$ and $\delta$, increasing the graph density leads to a higher $R$ for CPGR. For instance, when $m$ increases from 921 to 1,003, the relative $R$ rises from 1.03 to 1.25. In contrast, for a graph $n=128$, $m=16056$, density = 0.98, we observe a non-monotonic trend in the relative $R$ as $\delta$ increases. Initially, the ratio increases from 1.1 to 1.14 as $\delta$ goes from 0.5 to 0.7. However, further increasing $\delta$ to 1 results in a decrease to 1.1. This reduction is to the fact that a larger $\delta$ increases the size of the right partition $K_{q}$ of a $\delta$-clique (since $|K_{q}|=\hat{k}$ and $\hat{k}\propto\delta$), consequently decreasing the likelihood of finding common neighbors.
It is important to recognize that the optimal $R$ is influenced by the interplay between $n,m,$ and $\delta$. Identifying the specific combinations of these parameters that yield the highest reduction in the number of edges $R$ remains a subject for future investigation.